# The Impact of Imperfect XAI on Human-AI Decision-Making

### v3

@@ 6. Discussion
We investigate how imperfect XAI impacts humans’ decision-making when collaborating with an AI. More precisely, we assess how imperfect explanations affect humans’ reliance behavior and investigate the effects on the human-AI team performance. To answer RQ 1 and RQ 2, we assess the validity of our research model for two different types of explanations: natural language explanations and example-based explanations. Previous research emphasizes the need to consider imperfect AI when designing for human-AI collaboration (Kocielnik et al., 2019). With recent research looking into how humans and AI can achieve complementary team performance (Bansal et al., 2021), Schemmer et al. (2023) conceptualize the role of appropriate reliance in human-AI collaboration. We extend Schemmer et al. (2023)’s framework by adding another dimension: XAI advice. Given that an explanation can be incorrect even if the AI advice is correct, it is crucial to understand the impact of incorrect XAI advice on decision-making. Furthermore, it is necessary to understand the impact of imperfect XAI for different types of explanations. Below, we discuss how our contributions are situated in current literature and the implications for CSCW.
@@
In our study, we observe a significant moderation of humans’ level of expertise on the effect of explanations’ correctness on RAIR for both explanation modalities. However, we do not see this moderation for RSR. When humans are being provided wrong AI advice, their level of expertise does not moderate the impact of imperfect explanations on humans’ RSR. We do see a direct effect of the level of expertise on RSR in both explanation modalities. Additionally, the correctness of explanations impacts RSR negatively for example-based explanations. Overall, our work synthesizes how humans’ level of expertise impacts their reliance on AI when provided with imperfect explanations. Non-experts rely more on AI than experts, whereas experts rely more on their own initial decision. Especially for example-based explanations, imperfect XAI deceives experts’ self-reliance and experts’/non-experts’ relative AI reliance, inappropriately relying on the AI. Thus, this study sets a starting point to investigate the effect of imperfect XAI for different explanation modalities.
@@
Our findings show that imperfect explanations impact human-AI decision-making. We observe that experts reach complementary team performance when imperfect explanations are provided. This holds true for natural language and example-based explanations. While non-experts do not reach complementary team performance, our analyses reveal that their performance can be improved to be similar to that of the AI performance. Moreover, there is a difference between reaching complementary team performance when the correctness, or fidelity, of the explanation changes (see Figure 11 and Figure 12 in Section A.4). Previous research discusses the impact of explanations’ fidelity on humans’ reliance on AI and hypothesizes that fidelity has a positive impact on humans’ reliance behavior on AI (Hemmer et al., 2021). With our results, we confirm this hypothesis. Furthermore,  Papenmeier et al. (2019) observed that low-fidelity explanations (or incorrect explanations) impact user trust in AI when the global model performance is around $75\%$ accurate, which helps validate our findings. We also observe that the lack of expertise among non-experts impacts their task performance when shown incorrect explanations regardless of the AI advice being correct (Figure 12 in Section A.4). Similar to our findings, Nourani et al. (2020) observe that non-experts tend to over-rely on AI advice, attributing this to their inability to identify when the AI is incorrect because of their lack of expertise. These findings contribute to a more integrated understanding of the impact of human-AI decision-making on different user groups in the presence of imperfect XAI. For example, this can inform managers on how to assign tasks to humans with different levels of expertise and provide them with explanations in different modalities. It could also lead to organizations modifying their human-AI collaboration workflows. From informal conversations with the product team of an AI decision-support tool [^8] WildMe.org for biologists and conservationists to classify species and identify individuals from camera trap imagery, we learned that organizations using the tool have modified their workflow to incorporate “checks-and-balances”. For example, intro-level biologists will collaborate with the AI to match individuals and then request for a review of their “human-AI team” decision from a higher-up. In this unique human-human-AI collaboration scenario, the expert biologist could potentially correct situations when an intro-level biologist over-relies on the AI advice because of an incorrect explanation.
@@
Our findings can guide researchers and practitioners on how to assess and design for imperfect XAI in human-AI collaborations. Regardless of the explanation modality, it is important to understand how humans interact with imperfect XAI. Visual explanations, such as example-based explanations and saliency maps, have been shown in the past to be of high educational value to the end-user (e.g., (Kim et al., 2023b; Mac Aodha et al., 2018)), making it even more important to understand how to design for and mitigate imperfect XAI. This need is intensified with the role AI takes in organizational learning (Spitzer et al., 2023). Especially in the workplace, AI can facilitate knowledge transfer and support organizations in retaining and distributing expert knowledge (Spitzer et al., 2022; Jarrahi et al., 2023; Wilkens, 2020). Our findings can guide knowledge managers within organizations on how to make use of explanations for employees with different levels of domain knowledge. More precisely, knowledge managers should be aware of the impact of exposing humans with different levels of expertise to imperfect XAI. In addition, our findings contribute to a more integrated understanding of the impact that incorrect explanations can have on human-AI decision-making and inform different stakeholders in organizations. As non-experts are more affected by imperfect XAI — their performance drops more than experts’ performance when provided incorrect explanations in comparison to correct explanations. With this finding, knowledge managers can adjust their knowledge retention activities when training new employees; designers can adjust the development of human-AI collaboration systems to successfully facilitate explanations in decision-making and support humans in their work setting. Thus, we encourage practitioners designing human-AI collaboration systems to apply our findings to structure their design approach. This can aid organizations in laying out the strategic direction of human resource development by matching the use of explanations to humans’ prior knowledge. Overall, these findings shed light on the ongoing discussion in CSCW on how to make use of explanations within work settings.
@@ 7. Limitations & Future Work
Lack of Information to Properly Identify Birds. Expert birders usually rely on more information than just the visual characteristics of a bird when determining the bird species, especially for ambiguous cases. For example, the location and habitat in which the bird was spotted can be imperative to determine the exact bird species within a family. It’s unclear to what extent the lack of this information influenced our results. Future technical work could consider using this information to help build more transparent bird classification models.
@@
Simplistic Natural Language Explanations. The natural language explanations are very limited and simplistic, although this is a fault of the model that we are using to generate those natural language explanations (Hendricks et al., 2021). Compared to a field guide such as All About Birds (bir, 2023), these natural language explanations could be correct for several species given their lack of detail and their short length. This is potentially a side effect of the text descriptions in the CUB-200-2011 dataset being sourced from crowd workers instead of experts. We encourage technical researchers that are developing explanation methods that explain image classifications using natural language to pay careful attention to the text descriptions used for training the language model. For example, researchers could consider using text data from a field guide as their training data. We also encourage CSCW researchers to conduct more studies on how the explanation’s length and level of detail impact experts’ and non-experts’ appropriate reliance.
Different Explanations Convey Different Information. While we do not intend to directly compare the two explanation modalities throughout our analyses, they are discussed in terms of similarities and differences throughout the article. Our main intention is to investigate how our research model holds across different modalities of explanations. While several previous studies have compared multiple different types of explanation modalities qualitatively and quantitatively (e.g.,  (Kim et al., 2023b; Chen et al., 2023; Kim et al., 2023a; Du et al., 2022; Szymanski et al., 2021)), we encourage readers to avoid directly comparing the two explanations because they present different information. Previous research reveals that different explanation techniques can result in disagreements for the same dataset (Roy et al., 2022). For example, the natural language explanations from Hendricks et al. (2021) are feature-based, providing descriptions of features present in the image (Hendricks et al., 2021). However, the example-based explanations present three similar images, which is very different information from the natural language description of features. On top of that, the incorrectness in both explanation modalities is represented in different ways. While there are factual errors in natural language explanations that previous research addresses (e.g. hallucination effects of natural language models (Sovrano et al., 2023)), there are logical errors (e.g. inconsistencies) within example-based explanations. This opens avenues for future research to investigate how different human cognitive abilities (i.e. cognitive styles) impact the perception of these imperfect explanations in AI-assisted decision-making scenarios.
@@
AI Expertise versus Domain Expertise. Our analyses are based on the participant’s expertise in bird species identification. We do not ask participants about their knowledge of AI. Thus, we do not analyze their expertise related to AI. It is unclear to what extent the participant’s expertise with AI would impact the effect of imperfect XAI on appropriate reliance and task accuracy for bird species classification. Future work should consider looking at the effect that AI knowledge combined with domain expertise has on appropriate reliance taking into account an imperfect XAI.
@@ 8. Conclusion
This article sets out a research model to investigate the effect of imperfect XAI on human-AI decision-making. Thus far, human-computer interaction and CSCW literature lack to thoroughly scrutinize how explanations’ correctness affects humans’ decision-making and their reliance behavior on AI. Hence, through a human study with 136 participants, we empirically analyze humans’ decision-making and specifically assess whether their level of expertise and explanations’ assertiveness moderate the effect of imperfect XAI on appropriate reliance. Furthermore, we explore to what extent incorrect explanations deceive decision-makers’ reliance on AI. With our findings, we make several contributions: First, we propose a research model to investigate the moderation of assertiveness and humans’ level of expertise on imperfect XAI in decision-making tasks. We thereby extend the existing conceptualization of appropriate reliance by a new dimension of XAI advice. Second, through an empirical study, we reveal that imperfect explanations and participants’ level of expertise affect human-AI decision-making for two different explanation modalities. In addition, we show the effect on complementary team performance and provide guidance for future studies on how to investigate imperfect XAI in the context of human-AI decision-making. Third, we propose a novel metric called Deception of Reliance (DoR), which allows us to measure the impact of incorrect explanations on decision-makers’ reliance. Our results inform designers of human-AI collaboration systems and provide guidelines for their development. Fourth, we reveal which role the language tone in explanations plays and outline important dimensions that should be considered when designing for XAI advice.
@@ Acknowledgements.
We would like to thank Youwei Jiang for helping with the development of portions of the user interface used for the study. We would like to thank Hao-Fei Cheng and Haiyi Zhu for providing feedback on the initial study design. We also thank Max Schemmer and other lab members of the KSRI for discussing our study design and findings with us. We would like to thank Nari Johnson, Hayden Stec, Adel Sharif, Donny Bertucci, Alex Cabrera, Will Epperson, and Venkat Sivaraman and other members of the DIG Lab at CMU for participating in our initial pilot studies. We would like to thank all of the experienced birders that provided us with feedback on our study design and discussed the limitations of our study with us. Lastly, we would like to thank all of our connections that helped us recruit birders. Research reported in this publication was supported by the National Heart, Lung, and Blood Institute of the National Institutes of Health under award number R01HL164906. ChatGPT was utilized to generate LaTeX code for some of the tables in this paper.
@@ A.1. Appropriate Reliance with Imperfect XAI
$CSR_{IC}$
Correct self-reliance for the case where the AI gives incorrect advice and a correct explanation is one when the initial human decision is correct and the final decision is correct.
$OR_{IC}$
Over-reliance for the case where the AI gives incorrect advice and a correct explanation is one when the initial human decision is correct and the final decision is correct.
$CSR_{II}$
Correct self-reliance for the case where the AI gives incorrect advice and an incorrect explanation is one when the initial human decision is correct and the final decision is correct.
$OR_{II}$
Over-reliance for the case where the AI gives incorrect advice and an incorrect explanation is one when the initial human decision is correct and the final decision is correct.
$CAIR_{CC}$
Correct AI reliance for the case where the AI gives correct advice and a correct explanation is one when the initial human decision is incorrect and the final decision is correct.
$UR_{CC}$
Under-reliance for the case where the AI gives correct advice and a correct explanation is one when the initial human decision is incorrect and the final decision is correct.
$CAIR_{CI}$
Correct AI reliance for the case where the AI gives correct advice and an incorrect explanation is one when the initial human decision is incorrect and the final decision is correct.
$UR_{CI}$
Under-reliance for the case where the AI gives correct advice and an incorrect explanation is one when the initial human decision is incorrect and the final decision is correct.
@@ A.3. Moderation Analyses
coeff
ce
Z
p
LLCI
UCLI
const
1.26
.34
3.67
.00
.59
1.94
corr
.57
.53
1.07
.29
-.48
1.62
exp
-2.12
.33
-6.50
.00
-2.77
-1.48
Z1
-.46
.39
-1.16
.24
-1.23
.31
Z2
.00
.39
.00
1.00
-.76
.76
exp x corr
-1.00
.51
-1.96
.05
-1.99
.00
Z1 x corr
.46
.59
.77
.44
-.70
1.61
Z2 x corr
.19
.58
.33
.74
-.95
1.34
@@
coeff
ce
Z
p
LLCI
UCLI
const
.43
.32
1.35
.17
-.19
1.05
corr
1.02
.49
2.08
.04
.06
1.99
exp
-1.25
.31
-4.09
.00
-1.85
-.65
Z1
.26
.36
.73
.47
-.45
.98
Z2
.00
.37
.00
1.00
-.72
.72
exp x corr
-1.04
.47
-2.21
.03
-1.95
-.12
Z1 x corr
-.03
.54
-.06
.95
-1.08
1.02
Z2 x corr
-.16
.54
-.29
.77
-1.22
.90
@@
coeff
ce
Z
p
LLCI
UCLI
const
-17.16
592.16
-.03
.98
-1177.77
1143.45
corr
13.25
592.16
.02
.98
-1147.37
1173.86
exp
15.89
592.16
.03
.98
-1144.72
1176.50
Z1
.14
.52
-.26
.79
-.89
1.16
Z2
-.31
.56
-.56
.58
-1.41
.79
exp x corr
-13.33
592.16
-.02
.98
-1173.94
1147.29
Z1 x corr
-.28
.75
-.37
.71
-1.75
1.19
Z2 x corr
.90
.74
1.20
.23
-.56
2.36
@@
coeff
ce
Z
p
LLCI
UCLI
const
-3.60
.76
-4.74
.00
-5.09
-2.11
corr
-.44
1.29
-.34
.74
-2.97
2.10
exp
3.25
.74
4.39
.00
1.80
4.70
Z1
-.09
.43
-.22
.83
-.94
.75
Z2
.09
.43
.21
.83
-.75
.93
exp x corr
-.73
1.28
-.57
.57
-3.23
1.77
Z1 x corr
-.45
.75
-.60
.55
-1.92
1.02
Z2 x corr
-.43
.72
-.59
.55
-1.85
.99

### v4

@@ 6. Discussion
We investigate how imperfect XAI impacts humans’ decision-making when collaborating with an AI. More precisely, we assess how imperfect explanations affect humans’ reliance behavior and investigate the effects on the human-AI team performance. To answer RQ 1 and RQ 2, we assess the validity of our research model for two different types of explanations: natural language explanations and example-based explanations. Previous research emphasizes the need to consider imperfect AI when designing for human-AI collaboration (Kocielnik et al., 2019). With recent research looking into how humans and AI can achieve complementary team performance (Bansal et al., 2021), Schemmer et al. (2023b) conceptualize the role of appropriate reliance in human-AI collaboration. We extend Schemmer et al. (2023b)’s framework by adding another dimension: XAI advice. Given that an explanation can be incorrect even if the AI advice is correct, it is crucial to understand the impact of incorrect XAI advice on decision-making. Furthermore, it is necessary to understand the impact of imperfect XAI for different types of explanations. Below, we discuss how our contributions are situated in current literature and the implications for CSCW.
@@
In our study, we observe a significant moderation of humans’ level of expertise on the effect of explanations’ correctness on RAIR for both explanation modalities. However, we do not see this moderation for RSR. When humans are being provided wrong AI advice, their level of expertise does not moderate the impact of imperfect explanations on humans’ RSR. We identify a direct effect of the level of expertise on RSR in both explanation modalities. Additionally, the correctness of explanations impacts RSR negatively for example-based explanations. Overall, our work synthesizes how humans’ level of expertise impacts their reliance on AI when provided with imperfect explanations. Non-experts rely more on AI than experts, whereas experts rely more on their initial decisions. Especially for example-based explanations, imperfect XAI deceives experts’ self-reliance and experts’/non-experts’ relative AI reliance, fostering inappropriate reliance on the AI. Thus, this study sets a starting point to investigate the effect of imperfect XAI for different explanation modalities.
@@
Our findings show that imperfect explanations impact human-AI decision-making. We observe that experts reach complementary team performance when imperfect explanations are provided. This holds true for natural language and example-based explanations. While non-experts do not reach complementary team performance, our analyses reveal that their performance can be improved to be similar to that of the AI performance. Moreover, there is a difference between reaching complementary team performance when the correctness, or fidelity, of the explanation changes (see Figure 11 and Figure 12 in Section A.4). Previous research discusses the impact of explanations’ fidelity on humans’ reliance on AI and hypothesizes that fidelity has a positive impact on humans’ reliance behavior on AI (Hemmer et al., 2021). With our results, we confirm this hypothesis. Furthermore,  Papenmeier et al. (2019) observe that low-fidelity explanations (or incorrect explanations) impact user trust in AI when the global model performance is around $75\%$ accurate, which helps validate our findings. We also observe that the lack of expertise among non-experts impacts their task performance when shown incorrect explanations regardless of the AI advice being correct (Figure 12 in Section A.4). Similar to our findings, Nourani et al. (2020) observe that non-experts tend to over-rely on AI advice, attributing this to their inability to identify when the AI is incorrect because of their lack of expertise. These findings contribute to a more integrated understanding of the impact of human-AI decision-making on different user groups in the presence of imperfect XAI. For example, this can inform managers on how to assign tasks to humans with different levels of expertise and provide them with explanations in different modalities. It could also lead to organizations modifying their human-AI collaboration workflows. From informal conversations with the product team of an AI decision-support tool [^8] WildMe.org for biologists and conservationists to classify species and identify individuals from camera trap imagery, we learned that organizations using their tool have modified their workflow to incorporate “checks-and-balances”. For example, intro-level biologists will collaborate with the AI to match individuals and then request a review of their “human-AI team” decision from a higher-up. In this unique human-human-AI collaboration scenario, the expert biologist could potentially correct situations when an intro-level biologist over-relies on AI advice because of an incorrect explanation.
@@
Our findings can guide researchers and practitioners on how to assess and design for imperfect XAI in human-AI collaborations. Regardless of the explanation modality, it is important to understand how humans interact with imperfect XAI. Visual explanations, such as example-based explanations and saliency maps, have been shown in the past to be of high educational value to the end-user (e.g., (Kim et al., 2023b; Mac Aodha et al., 2018)), making it even more important to understand how to design for and mitigate imperfect XAI. This need is intensified with the role AI takes in organizational learning (Spitzer et al., 2023). Especially in the workplace, AI can facilitate knowledge transfer and support organizations in retaining and distributing expert knowledge (Spitzer et al., 2022; Jarrahi et al., 2023; Wilkens, 2020). Similarly, it is also crucial to understand how imperfect XAI affects the learning of novices through AI-based learning systems (Spitzer et al., 2023) or through collaboration with AI (Schemmer et al., 2023a). Our findings can guide knowledge managers within organizations on how to make use of explanations for employees with different levels of domain knowledge. More precisely, knowledge managers should be aware of the impact of exposing humans with different levels of expertise to imperfect XAI. In addition, our findings contribute to a more integrated understanding of the impact that incorrect explanations can have on human-AI decision-making and inform different stakeholders in organizations. As non-experts are more affected by imperfect XAI, their performance drops more than experts’ performance when incorrect explanations are provided in comparison to correct explanations. With this finding, knowledge managers can adjust their knowledge retention activities when training new employees; designers can adjust the development of human-AI collaboration systems to successfully facilitate explanations in decision-making and support humans in their work setting. Thus, we encourage practitioners designing human-AI collaboration systems to apply our findings to structure their design approach. This can aid organizations in laying out the strategic direction of human resource development by matching the use of explanations to humans’ prior knowledge. Overall, these findings shed light on the ongoing discussion in CSCW on how to make use of explanations within work settings.
@@ 7. Limitations & Future Work
Lack of Information to Properly Identify Birds. Expert birders usually rely on more information than just the visual characteristics of a bird when determining the bird species, especially for ambiguous cases. For example, the location and habitat in which the bird was spotted can be imperative to determine the exact bird species within a family. It’s unclear to what extent the lack of this information influences our results. Future technical work could consider using this information to help build more transparent bird classification models.
@@
Simplistic Natural Language Explanations. The natural language explanations are very limited and simplistic, although this is a fault of the model that we are using to generate those natural language explanations (Hendricks et al., 2021). Compared to a field guide such as All About Birds (bir, 2023), these natural language explanations could be correct for several species given their lack of detail and their short length. This is potentially a side effect of the text descriptions in the CUB-200-2011 dataset being sourced from crowd workers instead of experts. We encourage technical researchers who are developing explanation methods that explain image classifications using natural language to pay careful attention to the text descriptions used for training the language model. For example, researchers could consider using text data from a field guide as their training data. We also encourage CSCW researchers to conduct more studies on how the explanation’s length and level of detail impact experts’ and non-experts’ appropriate reliance.
Different Explanations Convey Different Information. While we do not intend to directly compare the two explanation modalities throughout our analyses, they are discussed in terms of similarities and differences throughout the article. Our main intention is to investigate how our research model holds across different modalities of explanations. While several previous studies compare multiple different types of explanation modalities qualitatively and quantitatively (e.g.,  (Kim et al., 2023b; Chen et al., 2023; Kim et al., 2023a; Du et al., 2022; Szymanski et al., 2021)), we encourage readers to avoid directly comparing the two explanations because they present different information. Previous research reveals that different explanation techniques can result in disagreements for the same dataset (Roy et al., 2022). For example, the natural language explanations from Hendricks et al. (2021) are feature-based, providing descriptions of features present in the image (Hendricks et al., 2021). However, the example-based explanations present three similar images, which is very different information from the natural language description of features. On top of that, the incorrectness in both explanation modalities is represented in different ways. While there are factual errors in natural language explanations that previous research addresses (e.g., hallucination effects of natural language models (Sovrano et al., 2023)), there are logical errors (e.g., inconsistencies) within example-based explanations. This opens avenues for future research to investigate how different human cognitive abilities (i.e., cognitive styles) impact the perception of these imperfect explanations in AI-assisted decision-making scenarios.
@@
AI Expertise versus Domain Expertise. Our analyses are based on the participant’s expertise in bird species identification. We do not ask participants about their knowledge of AI. Thus, we do not analyze their expertise related to AI. It is unclear to what extent the participant’s expertise with AI would impact the effect of imperfect XAI on appropriate reliance and task accuracy for bird species classification. Future work should consider looking at the effect that AI knowledge combined with domain expertise has on appropriate reliance, taking into account an imperfect XAI.
@@ 8. Conclusion
This article sets out a research model to investigate the effect of imperfect XAI on human-AI decision-making. Thus far, human-computer interaction and CSCW literature fail to thoroughly scrutinize how explanations’ correctness affects humans’ decision-making and their reliance behavior on AI. Hence, through a human study with 136 participants, we empirically analyze humans’ decision-making and specifically assess whether their level of expertise and explanations’ assertiveness moderate the effect of imperfect XAI on appropriate reliance. Furthermore, we explore to what extent incorrect explanations deceive decision-makers’ reliance on AI. With our findings, we make several contributions: First, we propose a research model to investigate the moderation of assertiveness and humans’ level of expertise on imperfect XAI in decision-making tasks. We thereby extend the existing conceptualization of appropriate reliance by a new dimension of XAI advice. Second, through an empirical study, we reveal that imperfect explanations and participants’ level of expertise affect human-AI decision-making for two different explanation modalities. In addition, we show the effect on complementary team performance and provide guidance for future studies on how to investigate imperfect XAI in the context of human-AI decision-making. Third, we propose a novel metric called Deception of Reliance (DoR), which allows us to measure the impact of incorrect explanations on decision-makers’ reliance. Our results inform designers of human-AI collaboration systems and provide guidelines for their development. Fourth, we reveal which role the language tone in explanations plays and outline important dimensions that should be considered when designing for XAI advice.
@@ Acknowledgements.
We would like to thank Youwei Jiang for helping with the development of portions of the user interface used for the study. We would like to thank Hao-Fei Cheng and Haiyi Zhu for providing feedback on the initial study design. We also thank Max Schemmer and other lab members of the KSRI for discussing our study design and findings with us. We would like to thank Nari Johnson, Hayden Stec, Adel Sharif, Donny Bertucci, Alex Cabrera, Will Epperson, Venkat Sivaraman, and other members of the DIG Lab at CMU for participating in our initial pilot studies. We would like to thank all of the experienced birders who provided us with feedback on our study design and discussed the limitations of our study with us. Lastly, we would like to thank all of our connections that helped us recruit birders. Research reported in this publication was supported by the National Heart, Lung, and Blood Institute of the National Institutes of Health under award number R01HL164906. ChatGPT was utilized to generate LaTeX code for some of the tables in this paper.
@@ A.1. Appropriate Reliance with Imperfect XAI
$CSR_{IC}$ Correct self-reliance for the case where the AI gives incorrect advice and a correct explanation is one when the initial human decision is correct and the final decision is correct.
$OR_{IC}$ Over-reliance for the case where the AI gives incorrect advice and a correct explanation is one when the initial human decision is correct and the final decision is correct.
$CSR_{II}$ Correct self-reliance for the case where the AI gives incorrect advice and an incorrect explanation is one when the initial human decision is correct and the final decision is correct.
$OR_{II}$ Over-reliance for the case where the AI gives incorrect advice and an incorrect explanation is one when the initial human decision is correct and the final decision is correct.
$CAIR_{CC}$ Correct AI reliance for the case where the AI gives correct advice and a correct explanation is one when the initial human decision is incorrect and the final decision is correct.
$UR_{CC}$ Under-reliance for the case where the AI gives correct advice and a correct explanation is one when the initial human decision is incorrect and the final decision is correct.
$CAIR_{CI}$ Correct AI reliance for the case where the AI gives correct advice and an incorrect explanation is one when the initial human decision is incorrect and the final decision is correct.
$UR_{CI}$ Under-reliance for the case where the AI gives correct advice and an incorrect explanation is one when the initial human decision is incorrect and the final decision is correct.
@@ A.3. Moderation Analyses
coeff ce Z p LLCI UCLI
const 1.26 .34 3.67 .00 .59 1.94
corr .57 .53 1.07 .29 -.48 1.62
exp -2.12 .33 -6.50 .00 -2.77 -1.48
Z1 -.46 .39 -1.16 .24 -1.23 .31
Z2 .00 .39 .00 1.00 -.76 .76
exp x corr -1.00 .51 -1.96 .05 -1.99 .00
Z1 x corr .46 .59 .77 .44 -.70 1.61
Z2 x corr .19 .58 .33 .74 -.95 1.34
@@
coeff ce Z p LLCI UCLI
const .43 .32 1.35 .17 -.19 1.05
corr 1.02 .49 2.08 .04 .06 1.99
exp -1.25 .31 -4.09 .00 -1.85 -.65
Z1 .26 .36 .73 .47 -.45 .98
Z2 .00 .37 .00 1.00 -.72 .72
exp x corr -1.04 .47 -2.21 .03 -1.95 -.12
Z1 x corr -.03 .54 -.06 .95 -1.08 1.02
Z2 x corr -.16 .54 -.29 .77 -1.22 .90
@@
coeff ce Z p LLCI UCLI
const -17.16 592.16 -.03 .98 -1177.77 1143.45
corr 13.25 592.16 .02 .98 -1147.37 1173.86
exp 15.89 592.16 .03 .98 -1144.72 1176.50
Z1 .14 .52 -.26 .79 -.89 1.16
Z2 -.31 .56 -.56 .58 -1.41 .79
exp x corr -13.33 592.16 -.02 .98 -1173.94 1147.29
Z1 x corr -.28 .75 -.37 .71 -1.75 1.19
Z2 x corr .90 .74 1.20 .23 -.56 2.36
@@
coeff ce Z p LLCI UCLI
const -3.60 .76 -4.74 .00 -5.09 -2.11
corr -.44 1.29 -.34 .74 -2.97 2.10
exp 3.25 .74 4.39 .00 1.80 4.70
Z1 -.09 .43 -.22 .83 -.94 .75
Z2 .09 .43 .21 .83 -.75 .93
exp x corr -.73 1.28 -.57 .57 -3.23 1.77
Z1 x corr -.45 .75 -.60 .55 -1.92 1.02
Z2 x corr -.43 .72 -.59 .55 -1.85 .99